The market value of Firm L's debt is $200,000 and its yield is 9%. The firm's equity has a market value of $300,000, its earnings are growing at a 5% rate, and its tax rate is 40%. A similar firm with no debt has a cost of equity of 12%. Under the MM extension with growth, what would Firm L's total value be if it had no debt?

a.

$417,000

b.

$437,850

c.

$377,286

d.

$358,421

e.

$397,143
Answer #1

total value of firm = 200000 + 300000 = 500000

total value = unlevered value + value tax shelter

=>

value tax shelter = 0.09 * 0.4/200000*(0.12-0.05)

= 102857

unlevered value = 500000 - 102857

= 397143

hence choose e)